Nick has a beautiful rectangular garden, which measures 4m by 3 m. He wants to create a uniform border of river rocks around three sides of his garden. If he wants the area of the border and the area of his garden to be equal, approximately how wide should the border be?
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

Total area - garden area = 12

Length =2x+4
width =x+3
(2x+4)(x+3)-12 =12
2x^2+6x+4x+12-12=12
2x^2+10x-12=0
divide by 2
x^2+5x-6=0
(x+6)(x-1)=0
x=-6 or x=1
X=-6 ignore
width =1 m

Answer by ikleyn(52780)

.

This problem admits and has another solution, too, if the three sides are chosen by another way.

The fact that the condition does not mention about this additional opportunity, is the FAULT of the problem's formulation.